OpenStudy (anonymous):

Which of the following expressions is the inverse of the function y equals quantity x minus 2 divided by 3? y = 3x + 2 y = 2x + 3 y equals quantity x minus 3 divided by 2 y equals quantity x plus 2 divided by 3

OpenStudy (onepieceftw):

Alright, first you need to swap x and y to solves for y. So you'll have x = x-2/3 Then you multiply both sides by 3 to get rid of the fraction. This gives us 3x = y-2 Now just add 2 to both sides and get y by itself. Now you have the function y = 3x+2

OpenStudy (anonymous):

swap x and y. now you have x=y-2/3. then you multiply both sides by three. that will get rid of the three on the right side. then you add two to both side and this will cancel the two out. then you get y=3x=2